OpenStudy (anonymous):

how do u factor k^2-16h^2

OpenStudy (anonymous):

You can write k^2-16h^2 as \(k^2 - (4h)^2 = (k-4h) \times (k+4h)\)

OpenStudy (anonymous):

ok